Keep getting it wrong because I need the answer in picofarads (pF) 1. 0/2 points Previous Answers PSE6 26.AE.01. Example 26.1 Parallel-Plate Capacitor Problem A parallel-plate capacitor has an
Answer #1

C=\frac{A\varepsilon _{0}}{d}=C=\frac{A\varepsilon _{0}}{d}=\frac{1.9*10^{-4}*8.854*10^{-12}}{1.26*10^{-3}}=1.335\ pF

C=\frac{A\varepsilon _{0}}{d}=\frac{1.9*10^{-4}*8.854*10^{-12}}{7.05*10^{-3}}=0.2386\ pF